What is the equilibirant vector of two vectors where one has a magnitude of 5cm pointin 60 deg N of E and the other has a magnitude of 3 cm poiting 30 deg S of W


1
Expert's answer
2021-03-02T18:05:27-0500
"\vec{R}=(5\cos{60},5\sin{60})+(-3\cos{30},-3\sin{30})\\=(-0.10,2.83)\ cm"

"R=\sqrt{0.1^2+2.83^2}=2.83\ cm\\\theta=\arctan{\frac{0.1}{2.83}}=1.98\degree\text{ W of N}"